aut = Urxe + 4tx 02x23, tro Их(Ort) = 1*, Их(3,t) =t to ИСХО)=1 tx, Oext3
Answer #1

244 = Uuxtdatxo LR<3 +70 i 7 using 9, 10,) and 13, +) --,+30 74 - (4,0) = 1+2 OLY23 Let u(x, t) = 12 t) + + x Substitute this

and the conditio 4443 the condition ulo) 1 WY, Y. becomes Uo) wie.) So wex,o=148o LX L3 (c) We most general solchon of equati

Here Ao - 23 Seite de (123 (16-) 1505 6 (1+x) cosme de [txisin ATX cos one t Ano a . iso 1 - 12 n is even nis odd thus, There